Please help me with these.
Add- 5b - 8b + 2b^2 and 3b^2 - 7b^3 + 5b.
Remove the parenthese and simplify if possible.
8d - ( -7c - 2d)
Subtract- 4x - 2x^2 + 4x^3 and 4x^3 + x - 3x^2.

Add using the vertical method.
5x^2 + 2x - 4, x^2 - 2x - 3, and 2x^2 -x - 3.
THANK YOU FOR YOUR HELP.

Answer by malakumar_kos@yahoo.com(315)


Add- 5b - 8b + 2b^2 and 3b^2 - 7b^3 + 5b.
Remove the parenthese and simplify if possible.
8d - ( -7c - 2d)
Subtract- 4x - 2x^2 + 4x^3 and 4x^3 + x - 3x^2.
Add using the vertical method.
5x^2 + 2x - 4, x^2 - 2x - 3, and 2x^2 -x - 3.

1) 5b-8b+2b^2+3b^2-7b^3+5b
5b+5b-8b+2b^2+3b^2-7b^3
b(5+5-8)+b^2(2+3)-7b^3
2b+5b^2-7b^3
2) 8d-(-7c-2d) = 8d+7c+2d = 8d+2d+7c = 10d+7c

3)4x-2x^2+4x^3-(4x^3+x-3x^2)
4x-2x^2+4x^3-4x^3-x+3x^2
3x^2-2x^2+4x-x = x^2+3x
4)5x^2+2x-4
x^2-2x-3
2x^2-x-3
adding the co-efficients of like terms ,we get
8x^2-x-10